OpenStudy (anonymous):

find the center of the circle x^2+y^2-8x-2y=19

OpenStudy (anonymous):

don't let me post answer

OpenStudy (anonymous):

\[x^2-8x+y^2-2y=19\]

OpenStudy (anonymous):

\[(x-4)^2+(y-1)^2=19+16+1\]

OpenStudy (anonymous):

\[(x-4)^2+(y-1)^2=36\]

OpenStudy (anonymous):

division both side by 36 you get \[\frac{(x-4)^2}{36}+\frac{(y-1)^2}{36}=1\]

OpenStudy (anonymous):

the center: ( 4,1)

OpenStudy (anonymous):

Wow. Thank you!

OpenStudy (anonymous):

uewc